La présentation est en train de télécharger. S'il vous plaît, attendez

La présentation est en train de télécharger. S'il vous plaît, attendez

SPS LS1 Planning meeting (N015) 19/06/13 David Mcfarlane

Présentations similaires


Présentation au sujet: "SPS LS1 Planning meeting (N015) 19/06/13 David Mcfarlane"— Transcription de la présentation:

1 SPS LS1 Planning meeting (N015) 19/06/13 David Mcfarlane

2 Topics Lifts and Monte-charges. Works affecting access. Magnet change campaign Planning. False floors open in the BAs Works at BA3 (Surface) E.C.Rs. Access rights and safety equipment Important Information. A.O.B

3 Lift Maintenance June/July

4 Works affecting access LocationWorksDateInterruption BA3 Safety works on Monte-Charge 24 th /25 th June Monte-Charge will be OUT OF SERVICE BA2 Safety works on Monte-Charge 26 th /27 th June Monte-Charge will be OUT OF SERVICE BA1Internet worksFriday the 28 th JuneNo Access into BA1 BA7 Safety works on Monte-Charge 3 rd /4 th July Monte-Charge will be OUT OF SERVICE BA6 Safety works on Monte-Charge 5 th /8 th July Monte-Charge will be OUT OF SERVICE

5 Magnet change campaign (week 24) MagnetLocationAccess viaDate MBB / BA5BA3/4/5 Monday 10/06/2013 (Morning) MBB / BA5BA3/4/5 Monday 10/06/2013 (Afternoon) MBA / BA4BA3/4 Tuesday 11/06/2013 (Morning) QD / BA5BA3/4/5 Tuesday 11/06/2013 (Afternoon) QD / BA5BA3/4/5 Wednesday 12/06/2013 (Morning) QD / BA4BA3/4 Wednesday 12/06/2013 (Afternoon) QF / BA4BA3/4 Thursday 13/06/2013 (Morning) QD / BA4BA3/4 Thursday 13/06/2013 (Afternoon)

6 Magnet change campaign (week 25) MagnetLocationAccess viaDate MBA / BA1BA6/1 Monday 17/06/2013 (Morning) MBB / BA4BA3/4 Monday 17/06/2013 (Afternoon) MBB / BA3 Tuesday 18/06/2013 (Morning) MBB / BA3 Tuesday 18/06/2013 (Afternoon) MBA / BA2BA3/2 Wednesday 19/06/2013 (Morning) MBB / BA2BA3/2 Wednesday 19/06/2013 (Afternoon) MBA / BA2BA3/2 Thursday 20/06/2013 (Morning) QNLD / TT20BA80/TT20 November 2013

7 EN UndergroundSur BA1 UndergroundSur BA2 UndergroundSur BA3 UndergroundSur BA4 UndergroundSur BA5 UndergroundSur BA6 TI8TI2 TRANSFER TUNNELS UndergroundSur BA7 UndergroundSur BA80/TCC Dec. Jan. Feb. Mar. Apr. May Jun. Jul. Aug. Sep. Oct. Nov. Dec. Jan. Feb. Mar. Apr. May Jun. Jul. Aug. Sep. Oct. Nov. Dec. Jan. Feb. Mar. Apr Dec. Jan. Feb. Mar. Apr. May Jun. Jul. Aug. Sep. Oct. Nov. Dec. Jan. Feb. Mar. Apr. May Jun. Jul. Aug. Sep. Oct. Nov. Dec. Jan. Feb. Mar. Apr. General maintenance Cold check out & Start-up DSO tests/HV tests EPC Commissioning tests Vacuum Ions to NA61 General Maintenance General Maintenance General Maintenance General maintenance General Maintenance UA9 works RF upgrade works EPC works Civil Engineering RF upgrade works RF Upgrade works Survey Horizontal Vertical smoothing Cables Irradiated Cables Cables Re-route Cabling Cables EL Studies Prep work in LSS1 Re-install equipment removed for cable works Survey SPS LS1 schedule Beam on Cool down/RP survey and Magnet tests Protons to the North Area Protons to LHC Cool down NO ACCESS Cool down Limited ACCESS Magnets Beam off : 16/02/13 Beam on: 13/10/14 Open Access : 27/03/13 Close Access : 27/06/14 Fibres CV works CV studies CV Controls upgrade Vacuum TT10 Consolidation 18 KV Septa works Kicker Septa works Kicker TT20 pipework Magnets 18 KV Civil Engineering Magnets Ring-main Vacuum Crab Cavities EPC works Cooling tower 863 Ring-main Cables Fibres 18 KV Civil Engineering in TT10 Fibres CV maintenance Cables Kicker Cables RAL Septa replacement

8 False floors lifted in Surface BAs BAsCabling pulling startCable pulling end BA1 (TS1- works) BA1 (Irradiated cable works) BA1 (Fibre Optics) 29/04/ /10/ /04/ /08/ /02/ /04/2013 BA206/01/201411/03/2014 BA3 BA3 (Fibre Optics) 27/03/ /04/ /05/ /05/2014 BA427/05/201312/07/2013 BA5 BA5 (Fibre Optics) 27/03/ /09/ /05/ /10/2013 BA6 BA6 (Fibre Optics) 30/09/ /5/ /10/ /05/2013 BA716/10/201307/11/2013 Below is a table showing the dates that EL will be pulling cables at each point of the SPS. During these periods large sections of the false floor in these surface buildings will be lifted. Any works in the BAs that involve lifting of the false floors MUST be reported to the Coordination team If you require access to racks or other equipment in the surface buildings during these periods you must check with planning first.

9 Works at BA3 Bonjour, Merci de trouver sur le doc ci-joint, le balisage à mettre en place pour les grutages Qui auront lieu du 24 au , entre le bat 870 et le bat 874. Merci pour votre collaboration. Serge Pelletier

10 10

11 New Cable Trench 2.5 to 3 weeks is needed to complete this work.

12 Engineering Change Requests Cher(e)s Collègues, Comme nous lavons évoqué à différentes reprises durant nos réunions de coordination, les modifications apportées à la machine doivent être signalées par lintermédiaire dune ECR (Engineering Change Request) qui véhicule lanalyse dimpact sur la machine. Un template standard a été développé principalement par les équipes EN-MEF-DC et DG-DI-DAT et approuvé la semaine dernière (LSC). Ce template est accessible de la façon suivante : Démarrer MS Word à File à New à PS-Complex ( ou à SPS ) à choisir le template que vous souhaitez Principes de base : Lutilisation de ce template couvre lensemble des modifications apportées aux équipements qui voient le faisceau (anneau principal et lignes de transfert), ainsi que les Services à alentours (CV, EL, HE, blindages, Cryo, …). Le remplacement dun équipement par un élément identique, la réparation dun équipement, les activités sur les infrastructures de génie civil ou sur les équipements de surface ne sont en principe pas soumis à cette procédure. LECR sera pourtant avantageusement utilisée dans certains cas (p.ex. interventions majeures) afin de permettre une coordination plus efficace et des interventions plus aisées des différentes équipes concernées. Un diagramme (flow chart) schématisant lutilisation du template et les interactions entre les différents intervenants est repris sous EDMS (version draft pour linstant – en cours de finalisation). Toutes les ECR sont centralisées par léquipe en charge de la Gestion de Configuration (EN-MEF-DC) pour TOUTES LES MACHINES. Chaque ECR, une fois pré-remplie par le responsable technique de lintervention, est donc envoyée à une adresse UNIQUE : Cette ECR est essentielle, et a pour objectifs de : Faire connaître au plus tôt le projet de modification de la machine ; Permettre à lensemble des personnes concernées démettre avis et commentaires, notamment au sujet de limpact de cette modification sur les équipements environnants ou le planning général ; Identifier les différents intervenants concernés par la modification ; Permettre un meilleur historique (traçabilité) des actions entreprises sur la machine ; Faciliter lorganisation et la coordination sur place de la modification ; Faciliter le suivi de la mise à jour des plans de layout concernés et la mise à jour dans CDD ; Faciliter le suivi de la mise à jour des banques de données machine (layout database ; HW baseline ; naming convention ; … ). Léquipe en charge de la Gestion de Configuration (EN-MEF-DC) fera alors circuler lECR via un processus dapprobation en EDMS. Ces approbations seront automatiquement diffusées en fonction de listes de diffusions copiées journellement depuis E-groups.E-groups Pour chaque machine, vous avez donc la possibilité de vous (dés)inscrire directement dans E-groups si vous souhaitez ou non participer au circuit dapprobation de ces ECR. Pour cela, les e-groups suivants ont été créés et sont en accès libre (CERN users) : ecr-sps Certains de ces e-groups contiennent déjà des noms, basés sur les listes précédentes, et il vous appartient de vous y (dés)inscrire. Nhésitez pas à contacter léquipe en charge de la Gestion de Configuration (EN-MEF-DC) via ou les coordinateurs techniques si vous avez des Davance je vous remercie pour votre collaboration.

13 Engineering Change Requests Templates and Guidelines for Engineering Change Requests: – EDMS: SPS ECR template (in DFS): – https://dfsweb.web.cern.ch/dfsweb/Services/DFS/DFSBrowser.aspx/A pplications/CERN/Workgroup%20Templates/SPS/ECR-Template- SPS.dotx https://dfsweb.web.cern.ch/dfsweb/Services/DFS/DFSBrowser.aspx/A pplications/CERN/Workgroup%20Templates/SPS/ECR-Template- SPS.dotx

14 Current ECRs for LS1 https://edms.cern.ch/nav/P:CERN :V0/P:CERN :V0/TAB3 1.Removal of AEWB Pick-up from position to (EDMS No ) RELEASED 2.Removal of a BPWA Pick-up from position to (EDMS No ) RELEASED 3.Installation of a new Wideband Kickers system in position (EDMS No ) RELEASED 4.Installation of a new HBW Damper Pick-up in place of BPWA & AEWB (EDMS No ) RELEASED 5.Replacement of 4 old BPA and 2 BPD by 6 BPCE in LSS1 & LSS2 (EDMS No ) RELEASED 6.Reinstallation of pump and gauges (Pirani + Penning) in (vacuum sector 150) (EDMS No ) RELEASED 7.Modification LSS1 sectorization (new VVSF + displacement VVSB 11772) (EDMS No v.2) RELEASED 8.Installation of 3 BCT in LSS5 for North Area protection during Ions operations in SPS. (EDMS No ) UNDER IMPLEMENTATION

15 Current ECRs for LS1 https://edms.cern.ch/nav/P:CERN :V0/P:CERN :V0/TAB3 9.Transfer of the Fast Beam Current Transformers (BCTFR.31451) for the SPS circulating beam from LSS3 to LSS5. (EDMS No ) RELEASED 10.SPS Beam Position Monitoring Prototyping Setup in LSS4. (EDMS No ) RELEASED 11.Installation de 4 moniteurs de position (2 BPH, 2 BPV), dédiés au système Damper. (EDMS No ) RELEASED 12.Installation d'un nouveau moniteur de position de type BPCR en LSS2 au (EDMS No ) RELEASED 13.Displacement of Door PPG2145 separating the LHC and the SPS in the TI2 tunnel. (EDMS No ) RELEASED 14.Interlocking of Injection Tunnel upstream areas with LHC powering. (EDMS No ) RELEASED 15.Move of BPH to BPH (EDMS No ) RELEASED 16.Optical Fibre Racks. (EDMS No ) RELEASED

16 Current ECRs for LS1 https://edms.cern.ch/nav/P:CERN :V0/P:CERN :V0/TAB3 17.Removal of old fixed TMP vacuum groups in (EDMS No ) RELEASED 18.BYBPM racks under the SPS dipole magnets as part of the Upgrade of Beam Position Monitors electronics (MOPOS). (EDMS No ) UNDER IMPLEMENTATION 19.Installation de Coffrets d'Alimentation en LSS1 et TT10. (EDMS No ) RELEASED 20.Installation d'un rack de jonction RX11695=TS1 au BA1, afin de limiter les longueurs des câbles à changer lors d'un futur remplacement. (EDMS No ) RELEASED 21.Installation de nouveaux chemins de câbles en LSS1 et LSS2. (EDMS No ) IN WORK

17 Access rights and safety equipment Everyone who enters the SPS MUST have the following: – Completed the correct courses. Safety course levels 1,2 and 3. Habilitation Electrique. Radiological Protection. Self Rescue Mask Training. – The correct access authorisation (via EDH). – A valid approved IMPACT. (with their name included on this IMPACT) – They must know the number of this IMPACT. – Safety shoes. – Biocell. – Electronic dosimetre. – Passive dosimetre. – Helmet plus Light. Please note that the medical service no longer contact you when your medical is about to expire!!!!!!! – You will receive an 2 months before informing you that some of your access rights will soon expire. – It is up to you to contact the medical service to arrange a medical.

18 Important information Useful Contact Numbers – SPS Coordination Team David MCFARLANE Franck BAIS – James RIDEWOOD – Control Room SPS CCC – TI CCC – CSA – RP Christophe Tromel – Angelito Herve – Nadine Conan (BA1) – Kurt Weiss (BA1) – Coordination Safety Emmanuel Paulat – John Etheridge

19 There are currently 38 IMPACTs that are LATE. If you have completed your work then you need to close the related IMPACT. There have been several thefts on the CERN sites recently (notably BA3,4 and 5) Please be vigilant and do not leave anything out over night. On Monday the 24 th June there will be a team doing 360 degree video scans of LSS1. This may cause some disruption to any teams working in this area. A.O.B.

20 SPS coordination website \\cernhomeJ.cern.ch\J\jocksoft\Public\LS1 Next meeting 03/07/ D20


Télécharger ppt "SPS LS1 Planning meeting (N015) 19/06/13 David Mcfarlane"

Présentations similaires


Annonces Google